a particle sarts from rest and moves  ina straight line with constant accelaration its velocity after n seconds is v the displacement of the body in the last two seconds is what

Grade:11

1 Answers

Himanshu
6 years ago
Displacement of body in `n` second =1/2v/n×n^2 =v×n/2Displacement of body in `n-2` second =1/2×v/n×(n-2)^2 =(n-2)^2×v/2nTherefore, acc to ques displacement of body in last 2 second= displacent in `n` second - displacenment in `n-2` second=v/2n[n^2-{n-2}^2] = v/2n(4n-4) =2v/n(n-1)